The net present value (NPV) of the Konka-Future Shop is R14 983 and the profitability index (PI) is 1,034. Approximately how much is the initial investment in the shop?

  • A. R14 983
  • B. R7 366
  • C. R440 676
  • D. R15 492
  • E. R14 490
1
Expert's answer
2021-11-21T15:22:54-0500

"Profitability Index = \frac{Net present value + Initial investment}{initial investment}"



"PI = \frac {NPV+IV}{IV}"


Finding initial investment


"1,034 = \frac {14983 + IV}{IV}"


"1,034IV = 14983+ IV"


"034IV = 14983"


"IV = \frac{14983}{034}"


Initial investment = "440 676" (c.)